# A systematic approach to Diophantine equations: open problems

### v1

@@ 1 Background and definitions
We start by recalling necessary definitions from [1]. A monomial in variables $x_{1},\dots,x_{n}$ with integer coefficient is any expression of the form $M=ax_{1}^{k_{1}}\dots x_{n}^{k_{n}}$, where $a$ is a non-zero integer and $k_{1},\dots,k_{n}$ are non-negative integers. Integer $d=\sum_{i=1}^{n}k_{i}$ is called the degree of monomial $M$. Two monomials $M=ax_{1}^{k_{1}}\dots x_{n}^{k_{n}}$ and $M^{\prime}=a^{\prime}x_{1}^{k^{\prime}_{1}}\dots x_{n}^{k^{\prime}_{n}}$ are called similar if $k_{i}=k^{\prime}_{i}$, $i=1,\dots,n$.
@@
It follows easily from the definition that, up to renaming the variables, there are only finitely many polynomial Diophantine equations of any given size. Hence, it is possible to order the equations by their size $H$, and solve them in this order.
@@ 2 Polynomial parametrization of all integer solutions
One of the basic questions one may ask about (1) is whether all its integer solutions can be parametrized.
@@ 3 Describing all integer/rational solutions
of size $H=32$ in integers.
@@
of size $H=26$, depending on whether you are satisfied with the description of integer solutions to (10) presented in [1]. Equation (11) is open without any doubts, because we do not even know whether it has infinitely many integer solutions that are not permutations of $(x,y,z)=(1-6t^{3},-6t^{2},1+6t^{3})$, $t\in{\mathbb{Z}}$.
@@
of size $H=26$. Equations (12) (a) and (12) (b) are equivalent, and so are the equations (13) (a) and (13) (b), see [1]. Whether equations (12) are open or not is debatable, because [1] contains some description of their solution sets, and everyone may decide for themselves whether to accept this as a “reasonable” description. Equations (13) and (14) are open without any doubts, because they have some obvious solutions, see [1], and it is open whether they have any other solutions.
@@
with $11/p+1/q+1/r<$ 1. An integer solution $(x,y,z)$ to (18) is called primitive if $\text{gcd}(x,y,z)=1$, and it is known that under condition $11/p+1/q+1/r<$ 1 the set of primitive solutions is finite. Table 4 presents the smallest equations of the form (18) for which the author of this document does not know the list of their primitive solutions.
@@ 4 Existence of arbitrary large integer solutions
Table 6: Three-variable equations of size $H\leq 64$ with open Problem 3.
@@ 6 Existence of integer solutions
of size $H=39$. The next-smallest open ones are the equations
@@ 7 The shortest open equations
$l$ Equation $l$ Equation $l$ Equation
$8$ $y(x^{3}-z^{2})=z$ $8$ $y(x^{3}-z^{2})=z+1$ $8$ $x^{2}y^{2}+xz^{2}+y-1=0$
$8$ $y(x^{3}-z^{2})=x$ $8$ $02x^{2}-xyz-y^{2}-1=$ 0 $8$ $x^{2}y^{2}+xz^{2}+y+1=0$
$8$ $y(x^{3}-z^{2})=x-1$ $8$ $02x^{2}-xyz+y^{2}+1=$ 0 $8$ $x^{3}y^{2}+z^{2}+y+1=0$
$8$ $y(x^{3}-z^{2})=x+1$ $8$ $y^{2}+x^{2}y+z^{2}x+1=0$ $8$ $y^{2}+x^{3}y+z^{2}+1=0$
Table 13: Equations of length $l\leq 8$ for which Problem 2 is open.
@@
Here, “possibly” depends on whether you are satisfied with the description of integer solutions to equations (39) presented in [1]. All these equations have length $l=7$.
@@
The shortest cubic equation for which Problem 6 is open is the equation
$06x^{2}z+z^{2}y+3y^{3}+1=$ 0 (40)
of length $10+2\log_{2}3\approx 13.2$. The next-shortest open cubic equations are the equation
$y^{2}+10xyz+x^{3}-x-2=0$
of length $11+\log_{2}5\approx 13.3$, and the equations
$y^{2}+7xyz+3x^{3}-2=0$
and
$17x^{3}+2y^{3}=3z^{2}+$ 1

### v2

# The smallest open polynomial Diophantine equations

@@ 1 Background and definitions
We start by recalling necessary definitions from [1, Section 1.1.1]. A monomial in variables $x_{1},\dots,x_{n}$ with integer coefficient is any expression of the form $M=ax_{1}^{k_{1}}\dots x_{n}^{k_{n}}$, where $a$ is a non-zero integer and $k_{1},\dots,k_{n}$ are non-negative integers. Integer $d=\sum_{i=1}^{n}k_{i}$ is called the degree of monomial $M$. Two monomials $M=ax_{1}^{k_{1}}\dots x_{n}^{k_{n}}$ and $M^{\prime}=a^{\prime}x_{1}^{k^{\prime}_{1}}\dots x_{n}^{k^{\prime}_{n}}$ are called similar if $k_{i}=k^{\prime}_{i}$, $i=1,\dots,n$.
@@
To compute the size $H(P)$ of any equation $P=0$, we just substitute $2$ in place of each variable, replace each coefficient by its absolute value, and evaluate. It follows easily from the definition that, up to renaming the variables, there are only finitely many polynomial Diophantine equations of any given size. Hence, it is possible to order the equations by their size $H$, and solve them in this order.
@@ 2 Polynomial parametrization of all integer solutions
One of the natural questions one may ask about (1) is whether all its integer solutions can be parametrized.
@@ 3 Describing all integer/rational solutions
of size $H=32$ in integers. We know that the corresponding cubic surfaces are unirational, see https://mathoverflow.net/questions/452461/, but currently do not know explicit parametrizations.
@@
of size $H=26$, depending on whether you are satisfied with the description of integer solutions to (10) presented in [1, Section 4.4.1]. Equation (11) is open without any doubts, because we do not even know whether it has infinitely many integer solutions that are not permutations of $(x,y,z)=(1-6t^{3},-6t^{2},1+6t^{3})$, $t\in{\mathbb{Z}}$.
@@
of size $H=26$. Equations (12) (a) and (12) (b) are equivalent, and so are the equations (13) (a) and (13) (b), see [1, Section 4.3.5]. Whether equations (12) are open or not is debatable, because [1, Section 4.3.5] contains some description of their solution sets, and everyone may decide for themselves whether to accept this as a “reasonable” description. Equations (13) and (14) are open without any doubts, because they have some obvious solutions, see [1, Section 4.3.5], and it is open whether they have any other solutions.
@@
with $11/p+1/q+1/r<$ 1. An integer solution $(x,y,z)$ to (18) is called primitive if $\text{gcd}(x,y,z)=1$, and it is known that under condition $11/p+1/q+1/r<$ 1 the set of primitive solutions is finite. Table 4 presents the smallest equations of the form (18) for which the author of this document does not know the complete list of their primitive solutions.
@@ 4 Existence of arbitrary large integer solutions
Table 6: Equations of size $H\leq 64$ with open Problem 3.
@@ 6 Existence of integer solutions
of size $H=39$. The next-smallest open symmetric ones are the equations
@@
$l$ Equation $l$ Equation $l$ Equation
$8$ $y(x^{3}-z^{2})=z$ $8$ $y(x^{3}-z^{2})=z+1$ $8$ $x^{3}y^{2}+z^{2}+y+1=0$
$8$ $y(x^{3}-z^{2})=x$ $8$ $y^{2}+x^{2}y+z^{2}x+1=0$ $8$ $y^{2}+x^{3}y+z^{2}+1=0$
$8$ $y(x^{3}-z^{2})=x-1$ $8$ $x^{2}y^{2}+xz^{2}+y-1=0$
$8$ $y(x^{3}-z^{2})=x+1$ $8$ $x^{2}y^{2}+xz^{2}+y+1=0$
Table 13: Equations of length $l\leq 8$ for which Problem 2 is open.
@@ 7 The shortest open equations
Here, “possibly” depends on whether you are satisfied with the description of integer solutions to equations (39) presented in [1, Section 4.3.5]. All these equations have length $l=7$.
@@
All equations in Table 15 have degree at least $4$. The shortest cubic equation for which Problem 6 is open is the equation
$17x^{3}+2y^{3}=3z^{2}+$ 1 (40)
@@
8 Changes in this document between versions
8.1 Changes between version 1 and version 2
Version 1 of this document corresponds exactly to equations left open in [1]. Below is the summary of updates between versions 1 and 2.
Equations $x^{4}+y^{4}+2z^{3}=0$ and $x^{5}+y^{4}+2z^{2}=0$ have been solved in the literature (see our forthcoming survey on generalized Fermat equations for details and references), hence these equations have been excluded from Table 4.
For the equations $02x^{2}-xyz-y^{2}-1=$ 0, $02x^{2}-xyz+y^{2}+1=$ 0 and $y^{2}+xyz+x^{3}+2=0$ Problem 2 has been solved111see https://mathoverflow.net/questions/449559/ and https://mathoverflow.net/questions/411958/. These equations have therefore been removed from Tables 5 and 13.
For the equations $y^{2}+xyz+x^{3}+2=0$, $y^{2}+xyz+x^{3}-3=0$, $y^{2}+xyz+x^{3}+4=0$, $y^{2}+xyz+x^{3}-x-2=0$, $y^{2}+xyz+y+x^{3}+2=0$ and $y^{2}+xyz+2x^{3}+1=0$ Problem 4 has been solved222see https://mathoverflow.net/questions/411958/ and https://mathoverflow.net/questions/466803/. These equations have been therefore removed from Tables 7 and 14.
For the equations $06x^{2}z+z^{2}y+3y^{3}+1=$ 0, $y^{2}+10xyz+x^{3}-x-2=0$ and $y^{2}+7xyz+3x^{3}-2=0$ Problem 6 has been solved333see https://mathoverflow.net/questions/466803/. Previously, these equations were the shortest cubic equations for which Problem 6 was open. These equations have now been removed from Section 7.